The temperature of an ideal gas remains constant while the absolute pressure changes from 100 kpa to 800 kpa. If initial volume is 100 liters, what is the final volume in liters?



1
Expert's answer
2022-03-24T16:36:53-0400

Answer

So final volume is

"V'=\frac{VP}{P'}\\=\frac{100*100}{800}\\=12.5liters"